The following costs have been estimated based on sales of 31,800 units: Total Annual Costs Percent That Is Variable Direct materials $ 379,200 100 % Direct labor 271,600 100 % Manufacturing overhead 268,000 50 % Selling and administrative 168,000 25 % What selling price (rounded to two decimal places) will yield a contribution margin of 40%? Multiple Choice $63.33 $47.50 $43.33 $37.00
Answer #1

Total Variable cost for 31800 units

= 379200+271600+(268000*50%)+(168000*25%)

= 826800

Variable cost per unit = 826800/31800 = 26

Let selling price per unit be x

(x-26)/x = 40%

x-26 = 0.40*x

x-0.40x = 26

0.60x = 26

x = 26/0.6

= 43.33